OpenStudy (anonymous):

will fan and medal but need help as soon as possible! Task 6—Polynomial Identities Pick a two-digit number greater than 25. Rewrite your two-digit number as a difference of two numbers. Show how to use the identity (x − y)2 = x2 − 2xy + y2 to square your number without using a calculator.

OpenStudy (loser66):

Pick 26 26 = 30-4 need show: \(26^2 = (30-4)^2\) by using the formula above. That is what they want you to do.

OpenStudy (anonymous):

so is that the answer?